Network Status

Current Status

All services are operational.

Recent Issues

12.05-1.20 pm, Friday, December 22nd - some services (including IdP) have been unavailable, requiring a restart of one of our servers, but al services should be available now. 

7.15 am, Wednesday, November 15th - we have noticed an issue that arose not long after 6 am that appears to be causing intermittent outages for some users.  The team is working on the issue and will resolve it ASAP.

9 am, Wednesday, November 15th - we have rebooted one of the servers which, as expected, appears to have cleared the problem.  Please let us know if there are any lingering issues. 

Past Issues

10.50am, Thursday, October 26th - we believe that most services have been back online but there has been some instability so we have just rebooted a subset of our servers to address that. Please advise us if there are any remaining issues.

9.10am, Thursday, October 26th - our apology as we have had an unexpected power issue at our Data Centre which may have briefly disrupted your service.  We believe this has been remedied now but please advise us if there are any remaining issues.

11.30am-1.45pm, Sunday, June 18th.  IdP services were down but all should be operational again now.  Please advise us if there are any lingering issues.

1 to 5 am, Wednesday, June 14th - our IdP was unavailable for a period overnight; all appears to be operational again but we are continuing checks.  Problem resolved

9.10pm, Friday, June 9th 9.10pm We are experiencing some intermittent upstream connectivity issues.  Problem resolved

12.45pm and 1.24pm, Wednesday, May 24th - we were aware of 2 problems briefly affecting the availability of services for some clients.  The issues were identified as related to upstream connectivity which the ISP concerned has subsequently confirmed.
Our services appear to all be operational but please advise us if you have any further difficulty and, in the meantime, we will continue to monitor the situation.  Problem resolved

8.45 to 10 am, Tuesday, May 9th - we are aware of a problem affecting the availability of services for some clients.  We are actively working on resolving the issue and will provide a further update as soon as possible.     Problem resolved.

5am to Noon, Sunday, March 26 - our systems have been acting unsteadily this morning and we have taken the opportunity to make some backend changes to help identify the recurring problems that we have experienced over the last few days,  Problem resolved.

12.40 to 2.30 pm, Saturday, March 25 - some of our services are again acting unsteadily but all services now appear to be fully operational.  Problem resolved.

5.20 to 6.10 pm, Thursday, March 23 - some of our services started acting unsteadily but all services now appear to be fully operational.  Problem resolved.

10.05 am, Wednesday, March 22 - the problem appears to have been resolved.  Please advise us at if you are still having an issue.  Problem resolved.

9.32 am, Wednesday, March 22 - there appears to be an issue and our engineers are looking into it urgently.

8.20am, Thursday, March 2, 2023 - SSO and hosted websites appear to be available again now.  If you have any services still unavailable (or underperforming), can you please advise us ASAP at Problem resolved - see Incident Report from Data Centre
7.30am, Thursday, March 2, 2023 - Most services are back online, unfortunately apart from SSO and hosted websites which we are still working on as quickly as possible.
6.30am, Thursday, March 2, 2023 - We understand that the Data Centre experienced a massive power surge and we are continuing to work on bringing all our services back online.
11.50pm, Wednesday, March 01, 2023 - A problem with our upstream provider is affecting some of our services.

11.20am, Wednesday, March 1st, 2023 - all services were down for approximately 30 minutes.   Problem resolved.

4pm, Tuesday, February 28th, 2023 - SSO services and some websites were down for approximately 1 hour.  Problem resolved.

4.30pm, Friday, November 11th, 2022 – the IdP was down for a little under 2 hours.  A reboot cleared the problem but the cause has yet to be determined.  Problem resolved.

8.30am, Tuesday, October 7th, 2021 – some clients have intermittent connectivity but it is proving very difficult to track down the cause.
At 10am this morning (Wednesday), we had sufficient data to escalate to our upstream provider who is helping us with the ongoing diagnostic work.  Problem resolved at approx 4pm on Thursday, October 7th.

8.30am, Thursday, Sept 30th, 2021 – the well-publicised LetsEncrypt issue that arose this morning affected a number of our services, especially single sign-on (SSO).  By midday, these appear to be (almost) all cleared.  Problem resolved.

1.30pm, Friday, Sept 3rd, 2021 – we have received reports of an inability to access the IdP and there appears to be an upstream connectivity issue involving Vocus/Orcon which we are looking into with our Data Centre provider.  This tuned out to be part of a much wider DDOS issue - see  Problem resolved.

1.35-2pm, Monday, May 17th, 2021 – we received reports of an SSO problem (ie users receiving a blank page) but a quiet reboot appears to have cleared the issue and we don’t expect a recurrence.  Problem resolved.

10.15am, Thursday, April 29th, 2021 – we have had a brief disruption but it appears that services are now back up and running, while we continue to investigate the cause of the issue.  Problem resolved.

5-8pm, Saturday, April 10th, 2021 – a disk issue unexpectedly disrupted a wide range of our services last evening; this has now been resolved and there is no likelihood of a recurrence.  Problem resolved.

6-8am, Thursday, January 28th, 2021 - a connectivity issue at our upstream provider (after their scheduled maintenance) caused all services to be inaccessible.  Problem resolved.

12.25pm, Monday, November 16th, 2020 - there was been a short, unscheduled outage due to a power issue caused by our Data Centre's upstream.  Problem resolved.

11.30am to 12.30pm, Friday, August 7th, 2020 - some services (including our IdP) will have been offline for a period but we believe everything should be operational again.  Problem resolved.

2.05 to 2.35pm, Monday, July 20th, 2020 - an IdP issue has now been resolved.  Problem resolved.

5.30 to 7.30am, Thursday, July 2nd, 2020 - an "international routing issue" at our upstream provider took down a number of our services.  Problem resolved.

Midnight to 12.30am, Thursday, June 11th, 2020 - scheduled maintenance by our Data Centre provider.

6 to 6.30am, May 26th, 2020 - scheduled maintenance successfully completed by our Data Centre provider.

7am to 9.05am, April 30th, 2020 - a routing issue at our upstream provider affected a number of our services, including our IdP.  Problem resolved.

6.30-11am, November 2nd, 2019 - some services (including the IdP) were down for a period.  Problem resolved.

10.30pm, July 16th to 10.45am, July 17th, 2019 - all services should now have resumed but, if anyone is having any ongoing issue, do please let us know ASAP.  Our apology for the unscheduled outage.  Problem resolved.

12.30 pm, March 28th, 2019 - the SSO service is now restored to full operation after being offline from earlier this morning.  Problem resolved.

11am, Jan 19th, 2019 - we appeared to have an issue briefly with DNS but this has now been resolved.  Problem resolved.

6am, Sept 27th, 2018 - there may be an issue with sending mail and the team is currently looking into it.  Problem resolved.

4pm, July 12th, 2018 - all services are operational (apart from Internet filtering for a small number of users, the cause of which should be resolved tomorrow by our upstream provider).  Problem resolved.
Noon, July 11th, 2018 - we have been bringing services back online progressively, with Internet filtering and outbound mail for some users still to be resolved; please let us know if there are any other lingering issues.
7am, July 11th, 2018 - we apologise for a disruption to some services, as we are working through an unexpected issue with our overnight upgrade.
9pm, July 10th, 2018 - as previously notified, we will be undertaking some important work this evening (as part of our infrastructure upgrade program) and some of our services will be unavailable for parts of a 2-3 hour period.

2.10pm, July 4th - there was apparently a brief issue with one of our hosted services but this has apparently been resolved now.  Problem resolved.

11pm, March 28th, 2018 - there was an unscheduled SSO outage for a period late last night/early morning which has now been resolved.  Problem resolved.
3pm, March 28th, 2018 - there appears to have been a short outage for our SSO service that has now been resolved.

11am, February 23rd, 2018 - there was an issue with outbound mail for at least some of our users and we believe this was resolved about 8pm (on the 23rd).  Problem resolved.

10am, January 5th, 2018 - as you may be aware, flaws have been identified in Intel (and other) chips which pose a potential security risk (see  We are monitoring the situation carefully and we will implement the requisite vendor updates as soon as they are available.  Please note that this could mean that at some point we need to take some of our services down at short notice but we will be looking to minimise any disruption.  If you have any queries, please contact us at  Problem resolved.

12.15pm, November 29th, 2017 - our IdP (and some other services) were affected by a brief disruption.  Problem resolved.

11.30am, November 24th, 2017 - our IdP (and some other services) were affected by an unscheduled outage for approximately 15 minutes but we believe that all services are operational again now.  Problem resolved.

5pm, October 11th, 2017 - to complete this part of our infrastructure upgrade program, we will be undertaking further work at this time; this may be for up to 2 hours and there is a modest possibility of an outage (for single sign-on and website hosting) for some or all of that time.  Problem resolved.

5.30pm, October 5th, 2017 - as part of our infrastructure upgrade program, we are undertaking some important work this evening; this may be for up to an hour and there is a modest possibility of an outage (for single sign-on and website hosting) for some or all of that time.  Problem resolved.

10am, August 28th, 2017 - it appears that some clients' recent DNS changes have been lost as a result of the problem on Friday but we are dealing with these as quickly as possible.  Problem resolved.

10.30am-1.30pm, August 25th, 2017 - some clients have been affected by slow (or no) performance of services this morning (in particular SSO, website hosting and DNS).  Our Support team has been working hard on resolving the issue and, subject to further checking, we believe it has now been resolved.  If you are continuing to have a difficulty, please contact us, preferably at  Problem resolved.

4pm, June 14, 2017 - there was a DNS issue that affected the availability of some of our services (including hosted websites) for a short period but this has now been resolved.  Problem resolved.

May 2nd-7th, 2017 - over the next few days, we are making some changes to our network as we finalise the decommissioning of some of Watchdog's old infrastructure; while we are not anticipating any disruption for end users, if you do notice any issues (especially with our mail or web filtering services) then please contact us so we can assist.  Problem resolved.

7.30pm, March 20th, 2017 - there will be some routine maintenance on our network between 7.30 and 8.30pm this evening and some services may experience a brief outage (<5 minutes)  Problem resolved.

5.50pm, March 9th, 2017 - some services may have had a short outage due to a hardware issue that has now been corrected.  Problem resolved.

12.30pm, March 3rd, 2017 - affected services are now progressively coming back online.

10.30am, March 3rd, 2017 - our initial resolution has apparently proved not to be successful so we are continuing work on the issue.

9am, March 3rd, 2017  - some services (eg SSO, some websites, DNS) are experiencing an outage which is being investigated.

6.30pm, February 28th, 2017  - there are no reports of any SSO client with an issue.  The Support team confirms that our service (and our underlying systems) have been working correctly all day (and in fact since early January) and all information available to us at the moment indicates a transitory upstream network issue.  We have made inquiries about this, including with our main upstream provider (Orcon/Vocus).

4.40pm, February 28th, 2017 - we have reports of an issue for some clients with the IdP.  Initial checks suggest it may be an upstream problem but we will continue to work on it.